OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

A. -1/7 B. -7/9 C. -8/7 D. -9 E. -4/3

OpenStudy (anonymous):

solve the two equations together by addition first multiple the 2nd eq. by 4 and then add the eqs

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

\[-9x+4y=8\] \[-12x-4y=16\]

OpenStudy (anonymous):

by addition -21x=24 x=-24/21=-8/7